A company is focus testing a new type of fruit drink . The focus group is 47% male. Of the responses, 40% of the males and 54% of the females said they would buy the fruit drink .What is the probability that a randomly selected person would buy the fruit drink ?
Answer by Theo(13342)

focus group is 47% male.
this means the focus group if 53% female.

40% of the males would be .4 * .47 = .188 of the group.
54% of the females would be .54 * .53 = .2862 of the group.

the total of the males that would buy the fruit and the females that would buy the fruit is .188 + .2862 = .4742 = 47.42% of the group.

assume the group contained 10,000 people.
47% male would be equal to 4700.
53% female would be equal to 5300.

40% of the males would be .4 * 4700 = 1880.
54% of the females would be .54 * 5300 = 2862.

the total members of the group that would buy the group is 1880 + 2862 = 4742.

4742 / 10,000 = .4742 = 47.42% of the group.

the percents and the numbers check out.
solution looks good.